OpenStudy (anonymous):

What is the value of the fourth term in a geometric sequence for which a1 = 15 and r = 1/3? Express your answer as a fraction.

OpenStudy (anonymous):

15/27

OpenStudy (lalaly):

15*(1/3)^3= 15/27=5/9

OpenStudy (anonymous):

thanks!